OpenStudy (anonymous):

4 3/4* divided by 3 2/3

OpenStudy (anonymous):

3/2

Parth (parthkohli):

= 4 3/4 * 3/11 = 19/4 * 3/11 = 57/44 =1 13/44

OpenStudy (anonymous):

4 3/4* divided by 3 2/3 (12/4) / (6/3) 3 / 2

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

3/2 is the exact?

Parth (parthkohli):

No, fontez is wrong. Please check your sources.

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

1 13/44 exact?

Parth (parthkohli):

4 3/4 = 19/4 not 12/4

OpenStudy (anonymous):

for steps: 4 3/4 = 3 ok? and 3 2/3 = 2 ok? result: 3/2

Parth (parthkohli):

4.75 is not equal to 12/4 that is 4

Parth (parthkohli):

Apply logic, fontez. see this too https://www.google.co.in/search?sourceid=chrome&ie=UTF-8&q=4+3%2F4+%2F+3+3%2F2

OpenStudy (anonymous):

Jyako90, 3/2 os correct

OpenStudy (anonymous):

3/2 is correct

Parth (parthkohli):

No.. it's not

OpenStudy (anonymous):

Im confused -_-

OpenStudy (anonymous):

if -> 4 3/4* divided by 3 2/3 ... is: 4(3/4) / 3(2/3) yes: 3/2

Parth (parthkohli):

4 3/4 = 4.75 not 4 as fontez suggests

OpenStudy (anonymous):

4 3/4 is 4(3/4) or 43/4 ????

Parth (parthkohli):

We're talking about mixed fractions not multiplication, fontez

OpenStudy (anonymous):

parthkohli? so whats right then?

Parth (parthkohli):

Look at the question. She says \[\Large 4 {3 \over 4}\] not 4(3/4)

OpenStudy (anonymous):

which is 19/4

OpenStudy (anonymous):

in my university is 3/2, bye. ;)

Parth (parthkohli):

Yes. fontez thought we're doing \[\Large 4({3 \over 4)}\]

Parth (parthkohli):

Haha I'm right :P

OpenStudy (anonymous):

lol =x

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

3/2?

OpenStudy (anonymous):

HEIII is 3/2, FINISH! :)

OpenStudy (anonymous):

& exact?

OpenStudy (anonymous):

ill let you guys know if i got it right or not lol

OpenStudy (anonymous):

question for your teacher ;)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

fontz whats the exact pls

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

estimate

OpenStudy (anonymous):

?

OpenStudy (anonymous):

my university: (4 3/4) / (3 2/3) equal (12/4) / (6/3) ok? in your school i dont no.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok ok but whats the estimate?

Parth (parthkohli):

He's just confusing you.

OpenStudy (anonymous):

pls explain

Parth (parthkohli):

What is 4 times 3/4?

OpenStudy (anonymous):

4/1* 3/4

OpenStudy (anonymous):

12/4

Parth (parthkohli):

Exactly. He thought that 4 3/4 meant 4 times 3/4

OpenStudy (anonymous):

lol wow...

OpenStudy (anonymous):

so is that the exact?

Parth (parthkohli):

You got it now? :P

Parth (parthkohli):

The answer is \[\LARGE 1 {13 \over 44}\]

Parth (parthkohli):

It's the exact answer

OpenStudy (anonymous):

exact

Parth (parthkohli):

If the estimate is asked, write 1. Exact is what I wrote \[1 {13 \over 44}\]

Parth (parthkohli):

YOU GOT IT NOW?

OpenStudy (anonymous):

ok so the estimate is? 14?

Parth (parthkohli):

The estimate is 1.

OpenStudy (anonymous):

kay thanks alot